Convert into a mole.
12 g of oxygen gas
Convert into a mole.
12 g of oxygen gas

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Given mass of oxygen gas = 12 g
Molar mass of oxygen gas (O2) = 32 g
Mole of oxygen gas 12 whole 12/32 = 0.375 mole

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Molar mass of Oxygen gas = 32 g
Given mass of Oxygen = 12 g
No. Of moles = Given mass/Molar mass
= 12/32 = 3/8
= 0.375 moles
 
Hope it helps!
Thank you ✌️